OpenStudy (anonymous):

Find an equation of a line that has the following: Slope=2, passes through (-5,-10) Slope= -1/3, passes through (-9,8)

OpenStudy (dtan5457):

Slope point formula= y-y1=m(x-x1) all you gotta do is plug them in, and simplify to whatever form your teacher asked you to do